# On second order conditions for singular optimal control of port-Hamiltonian systems

### v2

@@ 3.2 Optimality conditions for linear pH problems subject to control bounds
$\ddot{s}_{\mathcal{I}}$ $=p^{T}A^{2}G_{\mathcal{I}}$
$+$ $x^{T}\left[2QGYG^{T}QAG_{\mathcal{I}}-2A^{T}QGYG^{T}QG_{\mathcal{I}}+A^{T}A^{T}QGN_{\mathcal{I}}\right]$
@@
$\dot{x}_{1}$ $=$ $(\tilde{J}_{11}-\tilde{R}_{11})Q_{11}x_{1}+\tilde{G}_{1}u,$
$y$ $=$ $\tilde{G}_{1}^{T}Q_{11}x_{1},$
@@ 4 Conclusions and future work
Another classical application of optimal control is the computation of (feedback) controls $u=k(x)$ or output feedbacks $u=k(y)$ so that the closed loop system is asymptotically stable. To achieve this, one just chooses an infinite interval, i.e. $T=\infty$, and requires that for the closed loop solution $\lim_{t\to\infty}x(t)=0$. For this case, analogous results to the ones obtained in this work are expected to hold. It is also an interesting question what happens when the usual approach, of adding a quadratic penalty term $\alpha\sum_{j=1}^{m}u_{i}^{2}$ to the cost function is used, and one considers the convergence behavior for $\alpha\to 0$, see [15]. The Goh conditions will not appear in that context, since $\mathcal{H}_{u}$ depends on the control and, consequently, it is, in general, not differentiable in time.
We have studied problems without terminal constraints, which admit a unique multiplier that verifies normality, i.e., the multiplier associated to the integral cost can be normalized to $1$. In [3, 4] a framework for dealing with general terminal constraints of the form $\varphi_{i}(x(T))\leq 0,$ $\eta_{j}(x(T))=0,$ for $i=1,\dots,d_{\varphi},$ and $j=1,\dots,d_{\eta}$ was introduced. Also, [12] consider terminal constraints of the form $x(T)\in K$ under an additional hypothesis.
Using the techniques in [3, 4] one may as well discuss the derivation of second order sufficient conditions.
@@ Acknowledgment
Appendix A Appendix
In this Appendix we recall some of the general results for singular optimal control (1.7)-(1.9) given in Section 1. We further introduce the auxiliary notation
$f(x)\coloneqq f_{0}(x)+\sum_{i=1}^{m}f_{i}(x)u_{i},$
that simplifies the expressions in several upcoming calculations.
A.1 Optimality conditions and feedback formula
In this subsection we recall the optimality conditions for problem (1.7)-(1.9), with the assumption that the control $u$ takes values in an open set $\mathbb{U}\subseteq\mathbb{R}^{m}$.
The pre-Hamilton function (or un-maximized Hamilton function) for (1.7)–(1.9) reads (leaving off the arguments)
$\mathcal{H}=p^{T}f,$ (A.1)
where the adjoint vector function $p$ maps into $\mathbb{R}^{n}$, so that the partial derivative w.r.t $u_{i}$, for $i=1,\ldots,m$, is given by
$\mathcal{H}_{u_{i}}=\partial_{u_{i}}\mathcal{H}=p^{T}f_{i}.$ (A.2)
The adjoint equation in this case is given by
$-\dot{p}^{T}=\partial_{x}\mathcal{H}=p^{T}\partial_{x}f,\quad p(T)=\nabla\Psi(x(T)).$ (A.3)
By applying the Pontryagin Maximum Principle [34], it follows that the expression (A.2) vanishes along optimal trajectories, i.e.
$\mathcal{H}_{u_{i}}=0,\ i=1,\ldots,m.$ (A.4)
Since (A.4) does not depend explicitly on the control variable, one can differentiate with respect to time and obtains
$0=\frac{d}{dt}\mathcal{H}_{u_{i}}=\dot{p}^{T}f_{i}+p^{\top}f_{i}^{\prime}\dot{x}=p^{T}\big{(}[f_{0},f_{i}]+\sum_{j=1}^{m}u_{j}[f_{j},f_{i}]\big{)},$ (A.5)
where we used the state equation (1.8) and the adjoint equation (A.3) to replace $\dot{x}$ and $\dot{p},$ and where $f_{i}^{\prime}$ denotes the Jacobian, i.e., the derivative of $f_{i}$ with respect to $x$.
The following necessary condition is well-known, see e.g. [4, 12, 40].
Theorem A.1 (Goh condition).
For the optimal control problem (1.8), the Goh condition
$p^{T}[f_{i},f_{j}]=0,\quad\text{for }i,j=1,\dots,m$ (A.6)
holds along optimal trajectories.
Applying the Goh condition (A.6) to the expression $\frac{d}{dt}\mathcal{H}_{u_{i}}$ in (A.5) yields
$\frac{d}{dt}\mathcal{H}_{u_{i}}=p[f_{0},f_{i}],$ (A.7)
so that $\frac{d}{dt}\mathcal{H}_{u_{i}}$ is differentiable with respect to $t$. Differentiating (A.7) once more, one gets
$\frac{d^{2}}{dt^{2}}\mathcal{H}_{u_{i}}=p^{T}[f_{0},[f_{0},f_{i}]]+\sum_{j=1}^{m}u_{j}p^{T}[f_{j},[f_{0},f_{i}]].$
Thus, from $\frac{d^{2}}{dt^{2}}\mathcal{H}_{u_{i}}=0$ one obtains a linear system for $u$ given by (2.7), with the Hessian matrix $W=[w_{ij}]$ and the right hand side vector $d$, where for $i,j=1,\dots,m$,
$w_{ij}\coloneqq p[f_{j},[f_{0},f_{i}]],\qquad d_{i}\coloneqq p[f_{0},[f_{0},f_{i}]].$ (A.8)
This leads to a generalized Legendre-Clebsch condition, see e.g. [18, Theorem 6.2 and/or Corollary 6.3], [4], or [12].
Theorem A.2 (Generalized Legendre-Clebsch condition).
For the nonlinear control-affine problem (1.7)-(1.9), the associated Hessian matrix $W$ in (A.8) is negative semidefinite along optimal trajectories.
Observe that if $W$ is actually negative definite, then, from (2.7), one can express the optimal control in feedback form as
$u=-W^{-1}d.$ (A.9)
As already discussed in the introduction, the traditional technique of adding a regularizing term of the form $\alpha\sum_{i=1}^{m}u_{i}^{2}$ with positive $\alpha$ and then taking $\alpha\to 0$ does not work in this context, since (A.4) will depend on $u$ and then the expression (A.5) would not be possible. On the other hand, if one aims at obtaining a feedback formula for the controls on singular arcs by applying Pontryagin’s Maximum Principle to the modified quadratic cost and taking $\alpha\to 0,$ the expression contains $\alpha$ in the denominator and the limit may fail to exist, see, e.g. [15]. Despite this difficulty, it has been shown in the case of linear constant coefficients that this problem has some hidden regularity in a subspace [10].
Having reviewed the Goh conditions for the general unconstrained control-affine problem (1.7)-(1.9), in the next subsection we recall the optimality conditions for optimal control problems with control bounds.
A.2 Control-affine problems with control bounds
In this appendix we study problem (1.7)–(1.9) in the presence of bounds on the control as in (2.8).
In this case, the Goh optimality condition is restricted to the components that lie in the interior of the control set and one has the following optimality condition, see [4, 12].
Theorem A.3 (Goh condition under control constraints).
Consider problem (1.7)–(1.9) under the control bounds $u(t)\in\mathbb{U}$, for a.e. $t\in[0,T],$ with the set $\mathbb{U}$ as in (2.8). Then, on any open time interval over which the optimal control $u^{*}$ satisfies
it holds that (omitting arguments)
$p^{T}\,[f_{i},f_{j}]=0.$ (A.10)
Theorem A.4 (Generalized Legendre-Clebsch condition).
Under the hypotheses of Theorem A.3, let $(a,b)$ be an open interval on which
$\left\{\begin{array}[]{ll}\underline{u}_{i}<u^{*}_{i}(t)<\overline{u}_{i},&\text{for }i\in S,\\ u^{*}_{i}(t)=\underline{u}_{i},&\text{for }i\in\underline{B},\\ u^{*}_{i}(t)=\overline{u}_{i},&\text{for }i\in\overline{B}.\end{array}\right.$
Here $\{S,\underline{B},\overline{B}\}$ is a partition of $\{1,\dots,m\}$ into singular, lower-bang and upper-bang components of the control $u^{*}.$ Then, on $(a,b)$ and for $i\in S,$ it holds that
$p^{T}[f_{0},[f_{0},f_{i}]]+\sum_{j\in S}u_{j}^{*}p^{T}[f_{j},[f_{0},f_{i}]]$ (A.11)
$+\sum_{j\in\underline{B}}\underline{u}_{j}p^{T}[f_{j},[f_{0},f_{i}]]+\sum_{j\in\overline{B}}\overline{u}_{j}p^{T}[f_{j},[f_{0},f_{i}]]=0.$
Additionally, the $|S|\times|S|$-matrix $W_{S}(t)$ with entries $w_{ij}=p^{T}[f_{j},[f_{0},f_{i}]],$ for $i,j\in S$, is negative semidefinite.
Proof.
The expression (A.11) follows from calculations analogous to (A.2)–(A.5), (A.7)–(2.7) and Goh condition (A.10). The negative semidefiniteness of $W_{S}$ holds in view of [12, Theorem 4.1]. ∎
Remark A.1 (A feedback formula for the singular controls).
The matrix $W_{S}$ in Theorem A.4 is the submatrix of $W$ in (A.8) corresponding to the singular components of the control. Whenever $W_{S}$ is negative definite, by using expression (2.4), one can write the singular control in terms of the state variable $x$, the adjoint variable $p$ and the bang controls.

### v3

@@ 4.2 Optimality conditions for linear pH problems subject to control bounds
$\ddot{s}_{\mathcal{I}}=$ $\lambda^{T}A^{2}G_{\mathcal{I}}+x^{T}\left[2QGYG^{T}QAG_{\mathcal{I}}-2A^{T}QGYG^{T}QG_{\mathcal{I}}+A^{T}A^{T}QGN_{\mathcal{I}}\right]$
@@ 4.3 Linear descriptor systems
$\begin{split}\dot{x}_{1}&=(\tilde{J}_{11}-\tilde{R}_{11})Q_{11}x_{1}+\tilde{G}_{1}u,\\ y&=\tilde{G}_{1}^{T}Q_{11}x_{1},\end{split}$
@@ 5 Conclusions and future work
Another classical application of optimal control is the computation of (feedback) controls $u=k(x)$ or output feedbacks $u=k(y)$ so that the closed loop system is asymptotically stable. To achieve this, one just chooses an infinite interval, i.e. $T=\infty$, and requires that for the closed loop solution, $\lim_{t\to\infty}x(t)=0$. For this case, analogous results to the ones obtained in this work are expected to hold. It is also an interesting question what happens when the usual approach, of adding a quadratic penalty term $\alpha\sum_{j=1}^{m}u_{i}^{2}$ to the cost function is used, and one considers the convergence behavior for $\alpha\to 0$, see [JacGL70]. The Goh conditions will not appear in that context, since $\mathcal{H}_{u}$ depends on the control and, consequently, it is, in general, not differentiable in time.
We have studied problems without terminal constraints, which admit a unique multiplier that verifies normality, i.e., the multiplier associated to the integral cost can be normalized to $1$. In [Aronna2018, ABDL12] a framework for dealing with general terminal constraints of the form $\varphi_{i}(x(T))\leq 0,$ $\eta_{j}(x(T))=0,$ for $i=1,\dots,d_{\varphi},$ and $j=1,\dots,d_{\eta}$ was introduced. Also, [FraTon13] consider terminal constraints of the form $x(T)\in K$ under an additional hypothesis.
Using the techniques in [Aronna2018, ABDL12] one may as well discuss the derivation of second order sufficient conditions.